OpenStudy (anonymous):

Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (shamim):

is it sin 2x or sin^2x

OpenStudy (anonymous):

sin^2x

OpenStudy (shamim):

7sin^2x-14sin x+2+5=0 right?

OpenStudy (anonymous):

yes

OpenStudy (shamim):

7sin^2x-7sin x-7sin x+7=0

OpenStudy (shamim):

do u need any more help?

OpenStudy (anonymous):

how do I get the solutions? do I factor?

OpenStudy (shamim):

ya

OpenStudy (anonymous):

should my answer have pi in it?

OpenStudy (shamim):

7sinx(sinx-1)-7(sinx-1)=0

OpenStudy (shamim):

π radian=180 degree

OpenStudy (shamim):

u can convert degree to radian or radian to degree

OpenStudy (anonymous):

im very confused

OpenStudy (shamim):

ok. did u get my last step of ur math

OpenStudy (anonymous):

yes

OpenStudy (shamim):

(7sinx-7)(sinx-1)=0

OpenStudy (shamim):

7(sinx-1)(sinx-1)=0

OpenStudy (shamim):

(sinx-1)(sinx-1)=0

OpenStudy (anonymous):

when I did that I had another (sinx-1) 7(sinx-1)^3

OpenStudy (shamim):

now is it clear to u?

OpenStudy (anonymous):

no.. I don't know what to do with the sin x

OpenStudy (shamim):

did u get my last step

OpenStudy (anonymous):

(sinx-1)(sinx-1)=0

OpenStudy (shamim):

sinx-1=0

OpenStudy (anonymous):

how did you get rid of the other sinx-1

OpenStudy (shamim):

sinx=1

OpenStudy (shamim):

ok (sinx-1)^2=0 right?

OpenStudy (anonymous):

yes

OpenStudy (shamim):

now i wanna take square root of both side

OpenStudy (shamim):

sinx-1=0 right?

OpenStudy (anonymous):

ohhh gotcha

OpenStudy (anonymous):

So how do I know that that is the only answer?

OpenStudy (shamim):

sinx=1 sinx=sin 90 x=90 x=π/2

OpenStudy (shamim):

medal?

OpenStudy (anonymous):

I'm plugging those into my calculator and only 1 is working